OpenStudy (anonymous):

what is the acceleration of a body moving with constant acceleration covers the distance between two points 60m apart in 60sec? its velocity as it passes the second point is 15m/sec

OpenStudy (anonymous):

\[ \frac{ \Delta v}{ \Delta t} = a \]

OpenStudy (loser66):

@AllTehMaffs they didn't say it starts at rest, how can you know \(V_0 =0\)?

OpenStudy (anonymous):

you're right :( Sorry, I was bein' lazy. Then solve for u_0 in terms of acceleration and time \[ \frac{u-u_0}{t} = a \\ \ \\ u_0 = u - at \\ \ \\ s = (au-t) + \frac{1}{2}at^2 \]

OpenStudy (anonymous):

bah, I'm awful \[s = (u - at)t + \frac{1}{2}at^2 \]

OpenStudy (anonymous):

\[a = 2\frac{s - ut}{t^2}\]

OpenStudy (anonymous):

^_^

OpenStudy (anonymous):

Does that make sense @jontutz99 ?

OpenStudy (anonymous):

@AllTehMaffs what is u_O?

OpenStudy (anonymous):

but what formula should i use to get vo?

OpenStudy (anonymous):

The one you would normally use to find a constant acceleration \[ a=\frac{\Delta v}{t} = \frac{v-v_0}{t}\] \[ at = v - v_o\] \[v_0 = v - at\] we solve for it in terms of the velocity and time that we know ^^

OpenStudy (anonymous):

Then we can put that into the normal displacement equation \[ s = v_0t + \frac{1}{2}at^2 \] \[s = (v - at)t + \frac{1}{2}at^2\]

OpenStudy (anonymous):

ohh i get sir thanks :)

OpenStudy (anonymous):

Very welcome ^_^

OpenStudy (anonymous):

@AllTehMaffs sir you still online? i got a problem i get a - acceleration..

OpenStudy (anonymous):

That's okay. That just means it's slowing down ^_^

OpenStudy (anonymous):

sir should i use the formula s=(v−at)t+1/2at^2 to get the accel? or the other formula of a?

OpenStudy (anonymous):

cause i got a high Vo i got a 29.5 vo x.x i think im doing it wrong sorry

OpenStudy (anonymous):

hmm. That is fast1 ^_^ what did you get for a?

OpenStudy (anonymous):

And don't worry - getting things wrong is half the fun! ^^

OpenStudy (anonymous):

uhh.. if i use the s formula i get 0.466666666 if i use the a formula i get -0.466666666

OpenStudy (anonymous):

which means I made a mistake somewhere. I sec.. :P

OpenStudy (anonymous):

oh ok sir

OpenStudy (anonymous):

the acceleration formula is wrong, sorry about that. I messed up a sign. The top should be (vt - s). The acceleration then turns out to be positive, and the initial velocity is negative. darn +'s and -'s :P ^^

OpenStudy (anonymous):

Good intuition that the initial velocity was too big ^_^ I'm impressed!

OpenStudy (anonymous):

@AllTehMaffs i think we still got problem sir xD i got a high initial velocity .. the answer i got from the formula was 12.999999999 and the answer must be 1.67 so far x.x

OpenStudy (anonymous):

The correct answer is 1.67?

OpenStudy (anonymous):

yah 1.67 m/sec

OpenStudy (anonymous):

For the initial velocity?

OpenStudy (anonymous):

ahh nope for the acceleration oh sorry the answer i get for acceleration is 0.466666666 and the answer was suppose to be 1.67 sorry for that

OpenStudy (anonymous):

I'm really sorry that I've messed this up so many times for you :/ 1 sec...

OpenStudy (anonymous):

haha thats ok. important is you helping me :)

OpenStudy (anonymous):

And you're sure the time and displacement are t= 60s d = 60 m ?

OpenStudy (anonymous):

yah.. that was the given

OpenStudy (anonymous):

I have absolutely no idea how to get 5/3 as an answer. I'm sorry. There is some fundamental thing either missing in the problem or in my brain. If you use those numbers and just outright solve for v_o \[ d = \frac{v+v_0}{2}t\] \[ 2d/t = v+v_0\] \[2m/s - 15m/s = v_0 = -13m/s\] and that's what you get from the acceleration equation too. So it's not a miscalculation with any individual part of my setup. I have no idea. I think your answer is wrong. Or I'm crazy.

OpenStudy (anonymous):

xDD nvm sir thanks again for the help.. maybe ill try it nxt tme btw sir got any knowledge about parallel force?

OpenStudy (anonymous):

what's the problem? I'll try my best ^_^

OpenStudy (anonymous):

if a force of 86 N to the surface of 20 degree inclined plane will push a 120-N block up the plane at constant speed. what force parallel to the plane will push it down at constant speed?

OpenStudy (anonymous):

Can you draw a force diagram?

OpenStudy (anonymous):

uh.. sorry sir no fig here x.x

OpenStudy (anonymous):

So the first force is pushing completely horizontal?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh btw the answer here is 45 N

OpenStudy (anonymous):

My brain is not understanding what to do. I'm very, truly sorry. I need to sleep on it - can I get back to you tomorrow?

OpenStudy (anonymous):

yah sure ok thanks :)

OpenStudy (anonymous):

I apologize for my complete uselessness. G'night, and I'll try again tomorrow ^_^

OpenStudy (anonymous):

thats ok haha